OpenStudy (anonymous):

2 + (a + b) = (2 + a) + b is an example of which algebraic property? Select one: a. Distributive Property b. Associative Property of Addition c. Commutative Property of Addition d. Symmetric Property

Directrix (directrix):

Review the properties on the attached chart. http://tinyurl.com/og2arz5 @Jiujitsu4life

OpenStudy (anonymous):

Sorry about the long wait my friend, my computer is kinda slow as uaual. But let's get to it shall we: Associative Property = (a + b) + c = a + (b + c) Distributive Property = a(b + c) = ab + ac Communtative Property = a * b = b * a Symmetric Property: a = b the b = a Judging off of when you've seen here, i'm assuming you can find you're answer. I'm hoping i'm not too late. r